### v1

# On blow-ups of sets with finite fractional variation

Abstract.
Given $\alpha\in(0,1)$ and a set $E\subset\mathbb{R}^{N}$ with locally finite fractional $\alpha$-variation, we show that for almost every $x\in\mathbb{R}^{N}$ with respect to the $\alpha$-variation measure of $\mathbf{1}_{E}$, if $E$ admits a non-trivial tangent set at $x$ with locally finite integer perimeter, then $E$ also admits a tangent half-space oriented by the fractional unit normal of $E$ at $x$.
Key words and phrases:
Fractional gradient, fractional variation, tangent measure, blow-up, half-space
2020 Mathematics Subject Classification:
Primary 49Q15. Secondary 26A33, 28A75, 49Q20
Acknowledgements. The author is member of the Istituto Nazionale di Alta Matematica (INdAM), Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA), and has received funding from INdAM under the INdAM–GNAMPA Project 2025 Metodi variazionali per problemi dipendenti da operatori frazionari isotropi e anisotropi (grant agreement No. CUP_E5324001950001), and from the European Union – NextGenerationEU and the University of Padua under the 2023 STARS@UNIPD Starting Grant Project New Directions in Fractional Calculus – NewFrac (grant agreement No. CUP_C95F21009990001).
1. Introduction
1.1. Setting
Given $\alpha\in(0,1)$, the fractional $\alpha$-gradient of $u\in\operatorname{Lip}_{c}(\mathbb{R}^{N})$ is defined as
$\nabla^{\alpha}u(x)=c_{N,\alpha}\int_{\mathbb{R}^{N}}\frac{(u(y)-u(x))\,(y-x)}{|y-x|^{N+\alpha+1}}\,\mathrm{d}y,\quad x\in\mathbb{R}^{N},$ (1.1)
and the fractional $\alpha$-divergence of $\varphi\in\operatorname{Lip}_{c}(\mathbb{R}^{N};\mathbb{R}^{N})$ is analogously defined as
$\mathrm{div}^{\alpha}\varphi(x)=c_{N,\alpha}\int_{\mathbb{R}^{N}}\frac{(\varphi(y)-\varphi(x))\cdot(y-x)}{|y-x|^{N+\alpha+1}}\,\mathrm{d}y,\quad x\in\mathbb{R}^{N},$ (1.2)
where $c_{N,\alpha}>0$ is a suitable normalization constant. The operators (1.1) and (1.2) satisfy the integration-by-parts formula
$\int_{\mathbb{R}^{N}}u\,\mathrm{div}^{\alpha}\varphi\,\mathrm{d}x=-\int_{\mathbb{R}^{N}}\varphi\cdot\nabla^{\alpha}u\,\mathrm{d}x.$ (1.3)
For further details on the operators (1.1) and (1.2) and on the formula (1.3), we refer the reader to [silhavy20]. Building on the integration-by-parts formula (1.3), in our previous works [brue-et-al22, comi-stefani19, comi-et-al22, comi-stefani22, comi-stefani23a, comi-stefani23b, comi-stefani24a, comi-stefani24b], together with Giovanni E. Comi and several collaborators, we developed a new theory of distributional fractional Sobolev and $BV$ spaces.
1.2. Fractional variation
Given $p\in[1,\infty]$ and a non-empty open set $\Omega\subset\mathbb{R}^{N}$, we say that $u\in BV^{\alpha,p}_{\rm loc}(\Omega)$ if $u\in L^{p}(\mathbb{R}^{N})$ and
$|D^{\alpha}u|(A)=\sup\left\{\int_{\mathbb{R}^{N}}u\,\mathrm{div}^{\alpha}\varphi\,\mathrm{d}x:\varphi\in C^{\infty}_{c}(\mathbb{R}^{N};\mathbb{R}^{N}),\ \|\varphi\|_{L^{\infty}}\leq 1,\,\operatorname{supp}\varphi\subset A\right\}<\infty$
for every open set $A\Subset\Omega$. By Riesz’s Representation Theorem, $u\in BV^{\alpha,p}_{\rm loc}(\Omega)$ if and only if $u\in L^{p}(\mathbb{R}^{N})$ and there exists a locally finite (vector-valued) Radon measure $D^{\alpha}u\in\mathscr{M}_{\rm loc}(\Omega;\mathbb{R}^{N})$ such that (1.3) holds with $D^{\alpha}u$ in place of the fractional $\alpha$-gradient for every $\varphi\in C^{\infty}_{c}(\mathbb{R}^{N};\mathbb{R}^{N})$ such that $\operatorname{supp}\varphi\subset\Omega$. If $|D^{\alpha}u|(\Omega)<\infty$, then we write $u\in BV^{\alpha,p}(\Omega)$. Note that the subscript ‘loc’ in $BV^{\alpha,p}_{\rm loc}$ refers only to the local finiteness of the fractional variation measure, since $BV^{\alpha,p}_{\rm loc}$ functions are, by definition, in $L^{p}(\mathbb{R}^{N})$.
1.3. Blow-up Theorem
If $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$, then $|D^{\alpha}\mathbf{1}_{E}|\in\mathscr{M}_{\rm loc}(\mathbb{R}^{N})$ is the distributional fractional $\alpha$-perimeter measure of $E\subset\mathbb{R}^{N}$. In analogy with the classical theory [ambrosio-et-al00, maggi12], the fractional reduced $\alpha$-boundary $\mathscr{F}^{\alpha}E$ of $E$ is the set of points $x\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{E}|$ at which the (measure-theoretic) inner unit fractional normal exists, namely
$\nu^{\alpha}_{E}(x)=\lim_{r\to 0^{+}}\frac{D^{\alpha}\mathbf{1}_{E}(B_{r}(x))}{|D^{\alpha}\mathbf{1}_{E}|(B_{r}(x))}\in\mathbb{S}^{N-1}.$
The main feature of the fractional reduced boundary we are interested in here is its connection with the set of tangent (or blow-up) sets of $E$ at $x$, that is, the set $\operatorname{Tan}(E,x)$ of all limit points of $\left\{\frac{E-x}{r}:r>0\right\}$ in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ as $r\to 0^{+}$.
Theorem 1.1.
Let $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}^{\alpha}E$. Then, $\operatorname{Tan}(E,x)\neq\emptyset$ and any $F\in\operatorname{Tan}(E,x)$ is such that $\mathbf{1}_{F}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ with $\nu_{F}^{\alpha}=\nu_{E}^{\alpha}(x)$ for $|D^{\alpha}\mathbf{1}_{F}|$-a.e. $y\in\mathscr{F}^{\alpha}F$. Moreover, assuming $x=0$ and $\nu^{\alpha}_{E}(0)=\mathrm{e}_{N}$ without loss of generality, $F=\mathbb{R}^{N-1}\times M$ for some measurable set $M\subset\mathbb{R}$ such that:
  1. (i)
[…]
  2. (ii)
    $|M|,|M^{c}|\in\left\{0,\infty\right\}$;
  3. (iii)
    if $|M|=\infty$, then $\operatorname{ess}\sup M=\infty$;
  4. (iv)
    if $M\notin\{\emptyset,\mathbb{R}^{N}\}$ and $P(M)<\infty$, then $M=(m,\infty)$ for some $m\in\mathbb{R}$.
The first part of Theorem˜1.1 follows from [comi-stefani19]*Th. 5.8 and Prop. 5.9, while the second part is taken from [comi-stefani24b]*Th. 1.7. Theorem˜1.1 can be regarded as the fractional counterpart of De Giorgi’s Blow-up Theorem [maggi12]*Th. 15.5, which states that, if $\mathbf{1}_{E}\in BV_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}E$, the reduced boundary of $E$, then $\operatorname{Tan}(E,x)=\{H^{+}_{\nu_{E}(x)}(x)\}$, where
$H^{+}_{\nu_{E}(x)}(x)=\left\{y\in\mathbb{R}^{N}:(y-x)\cdot\nu_{E}(x)\geq 0\right\},$
and $\nu_{E}\colon\mathscr{F}E\to\mathbb{S}^{N-1}$ is the (measure-theoretic) inner unit normal of $E$. In fact, by [comi-stefani24b]*Th. 1.12(iii), if $\mathbf{1}_{E}\in BV_{\rm loc}(\mathbb{R}^{N})$, then $\mathscr{F}E\subset\mathscr{F}^{\alpha}E$ and $\nu_{E}^{\alpha}=\nu_{E}$ on $\mathscr{F}E$, showing that Theorem˜1.1 naturally extends the classical setting. However, the fractional reduced boundary $\mathscr{F}^{\alpha}E$ can be substantially larger than $\mathscr{F}E$, even when $E$ is very regular; see [comi-stefani24b]*Props. 1.13 and 1.15 for the cases of the half-space and the ball.
1.4. Main result
By combining Theorem˜1.1(iv) with well-known stability properties of the family of tangent sets (see [leonardi00]*Props. 2.1 and 2.2 for instance), we get that
$\text{$\mathbb{R}^{N-1}\times M\in\operatorname{Tan}(E,0)\setminus\{\emptyset,\mathbb{R}^{N}\}$ and $P(M)<\infty$}\implies H^{+}_{\mathrm{e}_{N}}(0)\in\operatorname{Tan}(E,0).$ (1.4)
This observation motivates the following question: does (1.4) still hold under the weaker assumption that $\mathbf{1}_{M}\in BV_{\rm loc}(\mathbb{R})$? The purpose of the present note is to provide an affirmative answer. Here and below, we write $BV_{\rm loc}^{\star}(\mathbb{R}^{N})=BV_{\rm loc}(\mathbb{R}^{N})\setminus\left\{0,1\right\}$ for brevity.
Theorem 1.2.
If $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$, then for $|D^{\alpha}\mathbf{1}_{E}|$-a.e. $x\in\mathscr{F}^{\alpha}E$ it holds that
$\operatorname{Tan}(E,x)\cap BV_{\rm loc}^{\star}(\mathbb{R}^{N})\neq\emptyset\implies H_{\nu_{E}(x)}^{+}(x)\in\operatorname{Tan}(E,x).$ (1.5)
Some comments are now in order. First, as shown by the aforementioned examples in [comi-stefani24b]*Props. 1.13 and 1.15, the implication (1.5) fails if $\operatorname{Tan}(E,x)=\left\{\emptyset\right\}$ or $\operatorname{Tan}(E,x)=\{\mathbb{R}^{N}\}$, so we must assume that a non-trivial blow-up exists. Second, Theorem˜1.2 is, at present, the closest result to De Giorgi’s Blow-up Theorem that we are able to achieve; namely, for $|D^{\alpha}\mathbf{1}_{E}|$-almost every $x\in\mathscr{F}^{\alpha}E$, if $E$ admits a non-trivial tangent set of locally finite integer perimeter at $x$, then some tangent of $E$ at $x$ is the half-space oriented by $\nu_{E}^{\alpha}(x)$. Third, we do not know whether this property holds for every $x\in\mathscr{F}^{\alpha}E$. In fact, it is not clear whether the proof of property (iv) in Theorem˜1.1 can be extended to the more general case $\mathbf{1}_{M}\in BV_{\rm loc}(\mathbb{R}^{N})$, thereby yielding Theorem˜1.2 for every $x\in\mathscr{F}^{\alpha}E$.
The statement of Theorem˜1.2—and its proof—was inspired by [ambrosio-et-al09]*Th. 1.2, where a similar result is obtained for sets of locally finite perimeter in Carnot groups. The idea underlying the proof of Theorem˜1.2, as well as that of the main result in [ambrosio-et-al09], is to exploit the principle—introduced by Preiss in [preiss87]*Th. 2.12 (see also [mattila95]*Th. 14.16)—that tangent measures to tangent measures are tangent measures (see Theorem˜2.1 below). Therefore, to obtain the conclusion of Theorem˜1.2, it suffices to ensure that at least one iterated tangent set of $E$ is a half-space, which is indeed the case whenever some non-trivial tangent set of $E$ has locally finite integer perimeter.
2. Preliminaries
2.1. Radon measure
Given a non-empty set $X\subset\mathbb{R}^{N}$, we let $\mathscr{M}(X)$ and $\mathscr{M}_{\rm loc}(X)$ be the spaces of finite and locally finite signed Radon measures on $X$, respectively.
[…]
$\lim_{k\to\infty}\int_{X}\varphi\,\mathrm{d}\mu_{k}=\int_{X}\varphi\,\mathrm{d}\mu\quad\text{for every }\varphi\in C_{c}(X).$ (2.1)
In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(X)$ as $k\to\infty$.
For vector-valued measures, we write $\mathscr{M}(X;\mathbb{R}^{m})$ and $\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$, with $m\in\mathbb{N}$. The notion of (local) weak convergence applies to $\mathbb{R}^{m}$-valued Radon measures by requiring (2.1) to hold componentwise. In this case, we write $\mu_{k}\xrightharpoonup{\star}\mu$ in $\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$ as $k\to\infty$. Accordingly, the total variation of $\mu\in\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$ on an open set $A\subset X$ is defined as
$|\mu|(A)=\sup\left\{\int_{X}\varphi\cdot\,\mathrm{d}\mu:\varphi\in C_{c}(X;\mathbb{R}^{m}),\ \operatorname{supp}\varphi\subset A,\ \|\varphi\|_{L^{\infty}}\leq 1\right\}.$
We recall that $|\mu|\in\mathscr{M}_{\rm loc}(X)$ for every $\mu\in\mathscr{M}_{\rm loc}(X;\mathbb{R}^{m})$; see [maggi12]*Lem. 4.17. For a detailed presentation of the theory of Radon measures, we refer to [ambrosio-et-al00, falconer97, maggi12, mattila95].
2.2. Tangent measures
Given $m\in\mathbb{N}$ and $\mu\in\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$, for every $x\in\mathbb{R}^{N}$ and $r>0$ we define $\mu_{x,r}\in\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$ by setting
$\mu_{x,r}(A)=\mu(x+rA)\quad\text{for every Borel set }A\subset\mathbb{R}^{N}.$
As customary (see [preiss87]*2.3(1), [mattila95]*Def. 14.1, and [ambrosio-et-al00]*Sec. 2.7), we say that $\nu\in\operatorname{Tan}(\mu,x)$ if and only if there exist sequences $(r_{k})_{k\in\mathbb{N}},(c_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ such that $r_{k}\to 0^{+}$ and
$c_{k}\,\mu_{x,r_{k}}\xrightharpoonup{\star}\nu\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$ as }k\to\infty.$
Moreover (see [falconer97]*Sec. 9.1), given $s\geq 0$, we say that $\nu\in\operatorname{Tan}_{s}(\mu,x)$ if there exists an infinitesimal subsequence $(r_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ such that
$r_{k}^{-s}\,\mu_{x,r_{k}}\to\nu\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$ as }k\to\infty.$
By definition, it is clear that $\operatorname{Tan}_{s}(\mu,x)\subset\operatorname{Tan}(\mu,x)$ for every $s\geq 0$, $\mu\in\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$, and $x\in\mathbb{R}^{N}$. For the proof of Theorem˜1.2, we will rely on the following crucial result.
Theorem 2.1.
Let $s\geq 0$, $m\in\mathbb{N}$ and $\mu\in\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{m})$. Then, for $|\mu|$-a.e. $x\in\mathbb{R}^{N}$, every $\nu\in\operatorname{Tan}_{s}(\mu,x)$ satisfies the following properties:
  1. (i)
    $\nu_{y,r}\in\operatorname{Tan}_{s}(\mu,x)$ for every $y\in\operatorname{supp}|\nu|$ and $r>0$;
  2. (ii)
    $\operatorname{Tan}_{s}(\nu,y)\subset\operatorname{Tan}_{s}(\mu,x)$ for every $y\in\operatorname{supp}|\nu|$.
For $m=1$ and $\mu\geq 0$, Theorem˜2.1 can be found in [falconer97]*Prop. 9.3 and, with $\operatorname{Tan}$ in place of $\operatorname{Tan}_{s}$, in [preiss87]*Th. 2.12 and [mattila95]*Th. 14.16. Moreover, Theorem˜2.1 corresponds to [ambrosio-et-al09]*Th. 6.4, stated on a Carnot group $\mathbb{G}$ in place of $\mathbb{R}^{N}$ and under the additional assumption that $\mu$ is asymptotically $s$-regular; that is,
$0<\liminf_{r\to 0^{+}}\frac{|\mu|(B_{r}(x))}{r^{s}}\leq\limsup_{r\to 0^{+}}\frac{|\mu|(B_{r}(x))}{r^{s}}<\infty\quad\text{for $|\mu|$-a.e.\ }x\in\mathbb{G}.$ (2.2)
Finally, Theorem˜2.1 also corresponds to [mattila05]*Prop. 2.15, with $m=1$, $\mathbb{R}^{N}$ replaced by a homogeneous locally compact metric group $\mathbf{G}$, and $\mu\geq 0$ such that
$\limsup_{r\to 0^{+}}\frac{\mu(B_{2r}(x))}{\mu(B_{r}(x))}<\infty\quad\text{for $\mu$-a.e.\ }x\in\mathbf{G}.$ (2.3)
The additional assumptions (2.2) and (2.3) are required to ensure the validity of the Differentiation Theorem, which does not generally hold in an arbitrary metric space.
The proof of Theorem˜2.1 follows almost verbatim the argument of [falconer97]*Prop. 9.3 (which, in turn, is nearly identical to that of [mattila95]*Th. 14.16), up to the minor modifications needed to treat the vector-valued case, as in [ambrosio-et-al09]*Th. 6.4. We therefore omit the details.
3. Proof of the main result
3.1. Properties of tangent sets
From Theorem˜2.1, we get the following result.
Corollary 3.1.
Let $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$. Then, for $|D^{\alpha}\mathbf{1}_{E}|$-a.e. $x\in\mathscr{F}^{\alpha}E$, every $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$ satisfies the following properties:
  1. (i)
    $\frac{F-y}{r}\in\operatorname{Tan}(E,x)$ for every $y\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{F}|$ and $r>0$;
  2. (ii)
    $\operatorname{Tan}(F,y)\subset\operatorname{Tan}(E,x)$ for every $y\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{F}|$.
In order to prove Corollary˜3.1, we need the following preliminary result.
Lemma 3.2.
If $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}^{\alpha}E$, then
$F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}\iff D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)\setminus\left\{0\right\}.$
Proof.
If $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$, then $\mathbf{1}_{F}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ by Theorem˜1.1. Moreover, we can find an infinitesimal sequence $(r_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ such that $\frac{E-x}{r_{k}}\to F$ in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ as $k\to\infty$. Therefore, by the scaling properties of the fractional $\alpha$-gradient (1.1) (see also [comi-stefani19]*Eq. (4.8)) and [comi-stefani24b]*Th. 1.6(i), we obtain
$r_{k}^{\alpha-N}(D^{\alpha}\mathbf{1}_{E})_{x,r_{k}}=D^{\alpha}\mathbf{1}_{\frac{E-x}{r_{k}}}\xrightharpoonup{\star}D^{\alpha}\mathbf{1}_{F}\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{N})$ as $k\to\infty$},$
showing that $D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)$. In addition, we must have $D^{\alpha}\mathbf{1}_{F}\neq 0$, since otherwise $\mathbf{1}_{F}$ would be constant by [comi-stefani24b]*Prop. 1.8, and thus $F\in\left\{\emptyset,\mathbb{R}^{N}\right\}$, a contradiction.
Conversely, if $D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)\setminus\left\{0\right\}$, then clearly $\mathbf{1}_{F}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$, and we can find an infinitesimal sequence $(r_{k})_{k\in\mathbb{N}}\subset(0,\infty)$ such that
$r_{k}^{\alpha-N}(D^{\alpha}\mathbf{1}_{E})_{x,r_{k}}\xrightharpoonup{\star}D^{\alpha}\mathbf{1}_{F}\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{N})$ as $k\to\infty$}.$
Thus, letting $E_{k}=(E-x)/r_{k}$ for every $k\in\mathbb{N}$, the sequence $(E_{k})_{k\in\mathbb{N}}$ satisfies $\mathbf{1}_{E_{k}}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ for every $k\in\mathbb{N}$, and
$D^{\alpha}\mathbf{1}_{E_{k}}=r_{k}^{\alpha-N}(D^{\alpha}\mathbf{1}_{E})_{x,r_{k}}\xrightharpoonup{\star}D^{\alpha}\mathbf{1}_{F}\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{N})$ as $k\to\infty$}.$
Since $x\in\mathscr{F}^{\alpha}E$, possibly passing to a subsequence (which we do not relabel), by Theorem˜1.1 there exists $G\in\operatorname{Tan}(E,x)$ such that $\mathbf{1}_{G}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $E_{k}\to G$ in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ as $k\to\infty$. Again by [comi-stefani24b]*Th. 1.6(i), possibly passing to a further subsequence (which we do not relabel), we also have
$D^{\alpha}\mathbf{1}_{E_{k}}\xrightharpoonup{\star}D^{\alpha}\mathbf{1}_{G}\quad\text{in $\mathscr{M}_{\rm loc}(\mathbb{R}^{N};\mathbb{R}^{N})$ as $k\to\infty$},$
from which it follows that $D^{\alpha}\mathbf{1}_{G}=D^{\alpha}\mathbf{1}_{F}$. By [comi-stefani24b]*Prop. 1.8, this implies that $\mathbf{1}_{G}-\mathbf{1}_{F}$ is constant, so that either $F=G$ or $F=\mathbb{R}^{N}\setminus G$. The latter possibility is ruled out, since otherwise $D^{\alpha}\mathbf{1}_{F}=-D^{\alpha}\mathbf{1}_{G}$ and thus $D^{\alpha}\mathbf{1}_{F}=0$, a contradiction. Hence $F=G$, and therefore $F\in\operatorname{Tan}(E,x)$. Moreover, $F\notin\{\emptyset,\mathbb{R}^{N}\}$, since $D^{\alpha}\mathbf{1}_{F}\neq 0$. ∎
Proof of Corollary˜3.1.
If $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$, then by Lemma˜3.2 we deduce that $D^{\alpha}\mathbf{1}_{F}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)\setminus\{0\}$. Therefore, by Theorem˜2.1(i), we get that $(D^{\alpha}\mathbf{1}_{F})_{y,r}\in\operatorname{Tan}_{N-\alpha}(D^{\alpha}\mathbf{1}_{E},x)$ for every $y\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{F}|$ and every $r>0$. Since $(D^{\alpha}\mathbf{1}_{F})_{y,r}=D^{\alpha}\mathbf{1}_{(F-y)/r}$ by the scaling properties of (1.1), again by Lemma˜3.2 this implies that $(F-y)/r\in\operatorname{Tan}(E,x)$ for every $y\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{F}|$ and $r>0$. Finally, since $\operatorname{Tan}(E,x)$ is closed with respect to convergence in $L^{1}_{\rm loc}(\mathbb{R}^{N})$ (see [leonardi00]*Prop. 2.2), we also obtain that $\operatorname{Tan}(F,y)\subset\operatorname{Tan}(E,x)$ for every $y\in\operatorname{supp}|D^{\alpha}\mathbf{1}_{F}|$, concluding the proof. ∎
3.2. Iterated tangent sets
For the proof of Theorem˜1.2, we rely on the notion of iterated tangent sets. Precisely, given $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$ and $x\in\mathscr{F}^{\alpha}E$, we define $\operatorname{Tan}^{1}(E,x)=\operatorname{Tan}(E,x)$ and
$\operatorname{Tan}^{k+1}(E,x)=\bigcup\left\{\operatorname{Tan}(F):F\in\operatorname{Tan}^{k}(E,x)\right\}$ (3.1)
for all $k\in\mathbb{N}$, where
$\operatorname{Tan}(E)=\bigcup\left\{\operatorname{Tan}(E,x):x\in\mathscr{F}^{\alpha}E\right\}.$
The proof of Theorem˜1.2 is based on the following result, which may be of independent interest (and rephrases [ambrosio-et-al09]*Th. 6.1 in the present setting).
Theorem 3.3.
If $\mathbf{1}_{E}\in BV^{\alpha,\infty}_{\rm loc}(\mathbb{R}^{N})$, then for $|D^{\alpha}\mathbf{1}_{E}|$-a.e. $x\in\mathscr{F}^{\alpha}E$ it holds that
$\bigcup_{k=2}^{\infty}\operatorname{Tan}^{k}(E,x)\subset\operatorname{Tan}(E,x).$
Proof.
Let $x\in\mathscr{F}^{\alpha}E$ be such that Corollary˜3.1(ii) holds. Then, for every $F\in\operatorname{Tan}(E,x)$ and every $y\in\operatorname{supp}\mathscr{F}^{\alpha}F$, we have $\operatorname{Tan}(F,y)\subset\operatorname{Tan}(E,x)$. By the definition in (3.1), we infer that $\operatorname{Tan}^{2}(E,x)\subset\operatorname{Tan}^{1}(E,x)$, and the conclusion follows by iteration. ∎
3.3. Proof of Theorem˜1.2
We can now prove our main result.
Proof of Theorem˜1.2.
Let $x\in\mathscr{F}^{\alpha}E$ be such that Theorem˜3.3 holds. By assumption, there exists $F\in\operatorname{Tan}(E,x)\setminus\{\emptyset,\mathbb{R}^{N}\}$ such that $\mathbf{1}_{F}\in BV_{\rm loc}(\mathbb{R}^{N})$. Hence, we have $\operatorname{Tan}(F,y)=\{H_{\nu_{F}(y)}^{+}(y)\}$ for every $y\in\mathscr{F}F$. Since $\mathscr{F}F\subset\mathscr{F}^{\alpha}F$ by [comi-stefani24b]*Th. 1.12(iii), from the definition in (3.1) we deduce that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}^{2}(E,x)$ for every $y\in\mathscr{F}F$. This, in turn, by Theorem˜3.3, implies that $H_{\nu_{F}(y)}^{+}(y)\in\operatorname{Tan}(E,x)$ for every $y\in\mathscr{F}F$. Combining [comi-stefani24b]*Prop. 1.13 with Theorem˜1.1, we then obtain that, for every $y\in\mathscr{F}F$,
$\nu_{F}(y)=\nu^{\alpha}_{H^{+}_{\nu_{F}(y)}(y)}(z)=\nu^{\alpha}_{E}(x)\quad\text{for a.e.}\ z\in\mathbb{R}^{N}.$
As a consequence, $\nu_{F}(y)=\nu^{\alpha}_{E}(x)$ for every $y\in\mathscr{F}F$, which implies that $F=H^{+}_{\nu^{\alpha}_{E}(x)}(x_{0})$ for some $x_{0}\in\mathbb{R}^{N}$ by [maggi12]*Prop. 15.15. Thus $H^{+}_{\nu^{\alpha}_{E}(x)}(x_{0})\in\operatorname{Tan}(E,x)$ for some $x_{0}\in\mathbb{R}^{N}$, and therefore $H^{+}_{\nu^{\alpha}_{E}(x)}(x)\in\operatorname{Tan}(E,x)$ as in (1.4), concluding the proof. ∎